One pipe fills a pool in 30 min
Another pipe fills a pool in 20 min
Another in 10 min
If all 3 were on how long till the pool is filled

Answer by mananth(16946)

one pipe fills 1/30 of pool in 1 minute
the second fills 1/20 of the pool in 1 minute
the third fills 1/10 of the pool in 1 minute
...
Together they will fill 1/10 + 1/20 + 1/30 of the pool in 1 minute
LCD = 60
(6+3+2)/ 60
=11/60 of the pool they fill in 1 minute
so they will take 60/11 minutes to fill the pool
60/11= 5 .45 minutes